# Some Constructions on Quantum Principal Bundles

### v4

@@ Abstract.
This paper works as an appendix of the paper titled Geometry of Associated Quantum Vector Bundles and the Quantum Gauge Group. Here, we are going to prove four statements in the theory of quantum principal bundles:
@@ 1. Universal Differential Envelope $\ast$–calculus in Differential Geometry
A $\mathcal{G}$–corepresentation $\delta^{V}$ is reducible if there exists a non–trivial subspace $L$ $(L\not=\{0\},V)$ such that $\delta^{V}(L)=L\otimes H$ and $\delta^{V}$ is unitary if viewed as an element of $B(V)\otimes H$ (with $B(V):=\{f:V\longrightarrow V\mid f\mbox{ is linear}\}$) is unitary. Of course, for the last definition, it is necessary an inner product $\langle$-|$\rangle$ - | - ⟩ on $V$. In [2], Woronowicz proved that for every finite–dimensional $\mathcal{G}$–corepresentation on $V$ there is an inner product $\langle$-|$\rangle$ - | - ⟩ (not necessarily unique) making the corepresentation unitary. Thus, henceforth we are going to consider that every finite–dimensional $\mathcal{G}$–corepresentation is unitary. In [2] one can find a proof of the following theorem.
@@
Let $\mathcal{T}$ be a complete set of mutually non–equivalent irreducible unitary (necessarily finite–dimensional) $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ (the trivial corepresentation on $\mathbb{C}$). For any $\delta^{V}$in $\mathcal{T}$ that coacts on $(V,\langle-|-\rangle)$,
@@
where $\{e_{i}\}^{n_{V}}_{i=1}$ is an ortonormal basis of $V$ and $\{g^{V}_{ij}\}^{n_{V}}_{i,j=1}$subseteq $H$. Then $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ is a linear basis of $H$, where the index $\delta^{V}$ runs on $\mathcal{T}$ and $i$, $j$ run from $1$ to $n_{V}=\mathrm{dim}_{\mathbb{C}}(V)$.
For every $\delta^{V}$in $\mathcal{T}$, the set $\{g^{V}_{ij}\}^{n_{V}}_{ij=1}$ satisfies
@@
where $[h]_{\mathcal{R}}$ denotes the equivalence class of the element $h$in $H$. The map $\pi$ has several useful properties, for example the restriction map $\pi|_{\mathrm{Ker}(\epsilon)}$ is surjective and
@@
for all $g$in $H$ [6]. It is worth mentioning that the canonical right action ${{}_{\Gamma}}\Phi$ of the $\ast$–FODC leaves $\mathfrak{qg}^{\#}$ invariant ([6]); so
@@
(11) $\mathrm{ad}\circ\pi=(\pi\otimes\mathrm{id}_{H})\circ\mathrm{Ad}.$
On the other hand there is a right $H$–module structure on $\mathfrak{qg}^{\#}$ given by
@@
for every $\theta=\pi(h)$in $\mathfrak{qg}^{\#}$; and it satisfies [6]
@@
for $g_{i}$, $h_{i}$in $H$. On the other hand, for a given $t=\vartheta_{1}\cdots\vartheta_{n}$in $\Gamma^{\wedge\,n}$ the linear map
@@
[…]
Let $(\Gamma,d)$ be a bicovariant $\ast$–FODC. Then the coproduct can be extended to a graded differential $\ast$–algebra morphism
@@
The counit and the coinverse can also be extended to $(\Gamma^{\wedge},d,\ast)$. These maps define a graded differential $\ast$–Hopf algebra structure
@@
on $\Gamma^{\wedge}$ which extends $H^{\infty}=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,d,\ast)$ [4]. In particular,
(18) $\epsilon:\Gamma^{\wedge}\longrightarrow\mathbb{C}$
given by $\epsilon|_{H}:=\epsilon$, $\epsilon|_{\Gamma^{\wedge k}}:=0$ for $k\geq 1$.
Now one can consider the right adjoint action of $\Gamma^{\wedge}$ by
@@
such that $\mathrm{Ad}(\vartheta)=(-1)^{\partial\vartheta^{(1)}\partial\vartheta^{(2)}}\vartheta^{(2)}\otimes S(\vartheta^{(1)})\vartheta^{(3)},$ where $\partial x$ denotes the grade of $x$ and $(\mathrm{id}_{\Gamma^{\wedge}}\otimes\Delta)\Delta(\vartheta)=(\Delta\otimes\mathrm{id}_{\Gamma^{\wedge}})\Delta(\vartheta)=\vartheta^{(1)}\otimes\vartheta^{(2)}\otimes\vartheta^{(3)}.$ Clearly, $\mathrm{Ad}$ extends the right adjoint coaction $\mathrm{Ad}$ of $H$.
@@
with $g$in $\mathcal{R}$ [4]. The space $\mathfrak{qg}^{\#\wedge}$ is a graded differential $\ast$–subalgebra ([4]) and it satisfies $\mathfrak{qg}^{\#\wedge}=\{\theta\in\Gamma^{\wedge}\mid\Phi_{\Gamma^{\wedge}}(\theta)=\mathbbm{1}\otimes\theta\}$. Also one can extend the right $H$–module structure of $\mathfrak{qg}^{\#}$ (see equation (12)) to $\mathfrak{qg}^{\#^{\wedge}}$ by means of
@@
Let $G$ be a compact matrix Lie group ($G\subseteq M_{k}(\mathbb{C}))$), and let $\mathcal{G}$ be its associated quantum group [2]. If $H^{\infty}=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,\ast,)$ is its dense $\ast$–Hopf algebra, then $H$ is the $\ast$–algebra generated by the smooth $\mathbb{C}$–valued functions
$w_{ij}:G$ $\longrightarrow\mathbb{C}$
@@
with $A=(a_{ij})$. The coproduct of an element $g$in $H$ is defined by
$\Delta(g):G\times G\longmapsto\mathbb{C},\qquad\Delta(g)(A,B)=g(AB)\quad\mbox{with }\quad A,\,B\,\in\,G.$
@@
$\epsilon:H\longrightarrow\mathbb{C},\qquad g\longmapsto g(e),$
where $e$in $G$ is the identity element; and
$S(g):G\longrightarrow\mathbb{C},\qquad A\longmapsto S(g)(A)=g(A^{-1})$
for all $g$in $H$.
Since $G$ is parallelizable, we have the identification $G\times\mathfrak{g}=TG$, where $\mathfrak{g}$ is the Lie algebra of $G$. This diffeomorphism can be given by
(24) $(A,v)\longmapsto(dL_{A})_{e}(v),$
where $(dL_{A})_{e}$ is the differential at $e$ of the left product on $G$ with $A$; and by
(25) $(A,v)\longmapsto(dR_{A})_{e}(v),$
where $(dR_{A})_{e}$ is the differential at $e$ of the right product on $G$ with $A$.
Under equation (24) we have $dg=g^{(1)}\otimes dg^{(2)}_{e}$ for all $g$in $H$, with $dh_{e}$ the differential of $h$in $H$ at $e$. By defining
@@
is a $\ast$–FODC over $\mathcal{G}$, where $d$ is the usual differential of smooth $\mathbb{C}$–valued functions. Moreover, it is left–covariant under the map $\Phi_{\Gamma}=\Delta\otimes\mathrm{id}_{\mathfrak{h}}$.
On the other hand, under equation (25), we have $dg=dg^{(1)}_{e}\otimes g^{(2)}$ for all $g$in $H$. Thus
$(\Gamma,d)=(\mathfrak{h}\otimes H,d)$
and it follows that $(\Gamma,d)$ is right–covariant by the map ${{}_{\Gamma}}\Phi=\mathrm{id}_{\mathfrak{h}}\otimes\Delta.$ Hence, $(\Gamma,d)$ is bicovariant. According to [4, 6], there exists a right $H$–ideal $\mathcal{R}\subseteq\mathrm{Ker}(\epsilon)$ such that
$\{\theta\in\Gamma\mid\Phi_{\Gamma}(\theta)=\mathbbm{1}\otimes\theta\}=\mathbbm{1}\otimes{\mathrm{Ker}(\epsilon)\over\mathcal{R}}={\mathrm{Ker}(\epsilon)\over\mathcal{R}}.$
@@
$\mathfrak{h}={\mathrm{Ker}(\epsilon)\over\mathcal{R}}=\mathfrak{qg}^{\#}.$
We claim that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon):=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,\mathrm{Ker}(\epsilon)\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$. In fact, consider the quantum germs map
@@
Then, by equation (9), we obtain $\pi(g)=S(g^{(1)})dg^{(2)}=dg_{e}$. If $\mathfrak{g}^{\#}_{\mathbb{C}}$ is the complexification of the dual space of $\mathfrak{g}$, it is well–known that for $X=\{f\in C^{\infty}_{\mathbb{C}}(G)\mid\epsilon(f)=0\}$ we have
@@
with $X^{2}=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,X\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$ [7]. Moreover, for every $f$in $X$,
@@
where $[f]_{X^{2}}$ denotes the equivalence class of $f$ in $X/X^{2}$ [7]. In this way, for every $g$in $\mathrm{Ker}(\epsilon)$ we get $dg_{e}=[g]_{\mathcal{R}}=[g]_{X^{2}}$. By this relation, $g$in $\mathcal{R}$ if and only if $g$in $X^{2}\cap\mathrm{Ker}(\epsilon)=\mathrm{Ker}^{2}(\epsilon),$ which proves that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon)$.
By definition, $\mathfrak{h}\subseteq\mathfrak{g}^{\#}_{\mathbb{C}}$. Since $\mathfrak{g}_{\mathbb{C}}\subseteq M_{k}(\mathbb{C})$, there exists a linear basis $\{v_{i}\}$ of $\mathfrak{g}_{\mathbb{C}}$ composed of linear combinations of the canonical linear basis $\{E_{ij}\}$ of $M_{k}(\mathbb{C})$ ($E_{ij}$ is the matrix with $1$ in the $(i,j)$ position and 00 elsewhere). Then $(dw_{ij})_{e}(v_{s})$ corresponds to the $(i,j)$ position of the matrix $v_{s}$. Consequently, there exist elements $\{(dg_{i})_{e}\}$, which are linear combinations of $(dw_{ij})_{e}$in $\mathfrak{h}$, such that $(dg_{i})_{e}(v_{j})=\delta_{ij}$ with $\delta_{ij}$ being the Kronecker delta. Thus, $\mathfrak{g}^{\#}_{\mathbb{C}}=\mathfrak{h}=\mathfrak{qg}^{\#}$ and we obtain
$(\Gamma,d)=(H\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d).$
This implies that $(\Gamma,d)$ is a $\ast$–subFODC of the $\ast$–FODC over $C^{\infty}_{\mathbb{C}}(G)$
$(C^{\infty}_{\mathbb{C}}(G)\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d)$
of $\mathbb{C}$–valued differential $1$–forms of $G$. Notice that it suffices to take convergent sequences $\{g_{i}\}^{\infty}_{i=1}$subseteq $H$ in $(\Gamma,d)$ to recover $(C^{\infty}_{\mathbb{C}}(G)\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d)$, although this lies beyond our purely geometric–algebraic approach. It is worth mentioning that for the specific form $\Gamma=H\otimes\mathfrak{g}^{\#}_{\mathbb{C}}$, the right covariance map ${{}_{\Gamma}}\Phi$ is given by [6]
@@
and hence the $\mathcal{G}$–corepresentation $\mathrm{ad}$ (see equation (10)) coincides with the dualization of the right adjoint action of $G$ on $\mathfrak{g}_{\mathbb{C}}$.
@@
$\pi(hg)=\pi(h)\epsilon(g)+\epsilon(h)\pi(g)$
and the right $H$–module structure on $\mathfrak{g}^{\#}_{\mathbb{C}}$, given by (see equation (12)) $\pi(h)\circ g=\pi(hg-\epsilon(h)g)$ simplifies to
$\pi(h)\circ g=\epsilon(g)\pi(h).$
@@
of $(\Gamma,d)$. If $g=a\,b$in $\mathrm{Ker}^{2}(\epsilon)$, then by equation (12), we obtain
@@
$\mathfrak{g}^{\#\,\wedge}_{\mathbb{C}}=\otimes^{\bullet}\mathfrak{g}^{\#}_{\mathbb{C}}/A^{\wedge}=\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}}.$
Thus (see equation (23))
$(\Gamma^{\wedge}=H\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast).$
@@
$(C^{\infty}_{\mathbb{C}}(G)\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast)$

### v5

@@ Abstract.
This paper works as an appendix of the paper titled Geometry of Associated Quantum Vector Bundles and the Quantum Gauge Group and for the paper titled Yang–Mills–Connes Theory and Quantum Principal SU(N)SU(N)–bundles. Here, we are going to prove four statements in the theory of quantum principal bundles:
@@ 1. Universal Differential Envelope $\ast$–calculus in Differential Geometry
A $\mathcal{G}$–corepresentation $\delta^{V}$ is reducible if there exists a non–trivial subspace $L$ $(L\not=\{0\},V)$ such that $\delta^{V}(L)=L\otimes H$ and $\delta^{V}$ is unitary if viewed as an element of $B(V)\otimes H$ (with $B(V):=\{f:V\longrightarrow V\mid f\mbox{ is linear}\}$) is unitary. Of course, for the last definition, it is necessary an inner product $\langle-|$-$\rangle$ on $V$. In [2], Woronowicz proved that for every finite–dimensional $\mathcal{G}$–corepresentation on $V$ there is an inner product $\langle-|$-$\rangle$ (not necessarily unique) making the corepresentation unitary. Thus, henceforth we are going to consider that every finite–dimensional $\mathcal{G}$–corepresentation is unitary. In [2] one can find a proof of the following theorem.
@@
Let $\mathcal{T}$ be a complete set of mutually non–equivalent irreducible unitary (necessarily finite–dimensional) $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$ $\in$ $\mathcal{T}$ (the trivial corepresentation on $\mathbb{C}$). For any $\delta^{V}$ $\in$ $\mathcal{T}$ that coacts on $(V,\langle-|-\rangle)$,
@@
where $\{e_{i}\}^{n_{V}}_{i=1}$ is an ortonormal basis of $V$ and $\{g^{V}_{ij}\}^{n_{V}}_{i,j=1}$ $\subseteq$ $H$. Then $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ is a linear basis of $H$, where the index $\delta^{V}$ runs on $\mathcal{T}$ and $i$, $j$ run from $1$ to $n_{V}=\mathrm{dim}_{\mathbb{C}}(V)$.
For every $\delta^{V}$ $\in$ $\mathcal{T}$, the set $\{g^{V}_{ij}\}^{n_{V}}_{ij=1}$ satisfies
@@
where $h]_{\mathcal{R}}$ denotes the equivalence class of the element $h$ $\in$ $H$. The map $\pi$ has several useful properties, for example the restriction map $\pi|_{\mathrm{Ker}(\epsilon)}$ is surjective and
@@
for all $g$ $\in$ $H$ [6]. It is worth mentioning that the canonical right action ${{}_{\Gamma}}\Phi$ of the $\ast$–FODC leaves $\mathfrak{qg}^{\#}$ invariant ([6]); so
@@
(11) $\mathrm{ad}\circ\pi=(\pi\otimes\mathrm{id}_{H})\circ\mathrm{Ad},$
where
(12) $\mathrm{Ad}:H\longrightarrow H\otimes H,\qquad g\longmapsto g^{(2)}\otimes S(g^{(1)})\otimes g^{(3)}$
is the (right) adjoint coaction of $H$.
On the other hand there is a right HH–module structure on $\mathfrak{qg}^{\#}$ given by
@@
for every $\theta=\pi(h)$ $\in$ $\mathfrak{qg}^{\#}$; and it satisfies [6]
@@
for $g_{i}$, $h_{i}$ $\in$ $H$. On the other hand, for a given $t=\vartheta_{1}\cdots\vartheta_{n}$ $\in$ $\Gamma^{\wedge\,n}$ the linear map
@@
given by $d(t)=d(\vartheta_{1}\cdots\vartheta_{n})=\displaystyle\sum^{n}_{j=1}(-1)^{j-1}\vartheta_{1}\cdots\vartheta_{j-1}\cdot d\vartheta_{j}\cdot\vartheta_{j+1}\cdots\vartheta_{n}$ $\in$ $\Gamma^{\wedge\,n+1}$ is well–defined, satisfies the graded Leibniz rule and $d^{2}=0$ [4, 6]. In other words, $d$ is natural extension in $\Gamma^{\wedge}$ of the differential of $\Gamma$. In this way $(\Gamma^{\wedge},d,\ast)$ is a graded differential $\ast$–algebra generated by $\Gamma^{\wedge\,0}=H$ (the degree 0 elements) [4, 6].
Let $(\Gamma,d)$ be a bicovariant $\ast$–FODC. Then, the coproduct can be extended to a graded differential $\ast$–algebra morphism
@@
In particular, in accordance with [4, 6], we have
(18) $\Delta(\theta)=\mathbbm{1}\otimes\theta+\mathrm{ad}(\theta).$
for all $\theta$ $\in$ $\mathfrak{qg}^{\#}$.
The counit and the coinverse can also be extended. In fact, consider the linear map
(19) $\epsilon:\Gamma^{\wedge}\longrightarrow\mathbb{C}$
defined by $\epsilon|_{H}:=\epsilon$ and $\epsilon|_{\Gamma^{\wedge k}}:=0$ for $k\geq 1$.
On the other hand, for any $g$ $\in$ $\mathrm{Ker}(\epsilon)$, define
$S^{1}(g):=-\pi(g^{(2)})S(g^{(3)})S(S(g^{(1)})).$
Since $\mathrm{Ad}(\mathcal{R})\subseteq\mathcal{R}\otimes H$, we obtain $S^{1}(g)=0$ for all $g$ $\in$ $\mathcal{R}$. Hence, there exists a well–defined linear map $S^{1}:\mathfrak{qg}^{\#}\longrightarrow\Gamma.$ In accordance with [4], we can extend $S^{1}$ to the whole $\Gamma$ in such a way that the following equalities hold
$S^{1}(h\,\pi(g))=S^{1}(\pi(g))\,S(h)\quad\mbox{ and }\quad S^{1}(h\,dg)=d(S(g))\,S(h)$
for all $g$, $h$ $\in$ $H$. According to [4, 6], we extend $S^{0}:=S$ and $S^{1}$ to a graded antimultiplicative linear map
(20) $S:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}$
which commutes with the differential ([4]). These maps define a graded differential $\ast$–Hopf algebra structure
@@
on $\Gamma^{\wedge}$ which extends $H^{\infty}=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,d,\ast)$ [4].
Now it is possible to consider the right adjoint coaction of $\Gamma^{\wedge}$ by taking
@@
such that
$\mathrm{Ad}(t)=(-1)^{\partial t^{(1)}\partial t^{(2)}}t^{(2)}\otimes S(t^{(1)})t^{(3)},$
where $\partial x$ denotes the grade of $x$ and $(\mathrm{id}_{\Gamma^{\wedge}}\otimes\Delta)\Delta(t)=(\Delta\otimes\mathrm{id}_{\Gamma^{\wedge}})\Delta(t)=t^{(1)}\otimes t^{(2)}\otimes t^{(3)}.$ Clearly, $\mathrm{Ad}$ extends the right adjoint coaction $\mathrm{Ad}$ of $H$.
@@
with $g$ $\in$ $\mathcal{R}$ [4]. The space $\mathfrak{qg}^{\#\wedge}$ is a graded differential $\ast$–subalgebra ([4]) and it satisfies $\mathfrak{qg}^{\#\wedge}=\{\theta\in\Gamma^{\wedge}\mid\Phi_{\Gamma^{\wedge}}(\theta)=\mathbbm{1}\otimes\theta\}$. Also one can extend the right HH–module structure of $\mathfrak{qg}^{\#}$ (see equation (13)) to $\mathfrak{qg}^{\#^{\wedge}}$ by means of
@@
Let $G$ be a compact matrix Lie group ($G\subseteq M_{k}(\mathbb{C}))$), and let $\mathcal{G}$ be its associated quantum group [2]. If
(27) $H^{\infty}=(H,\cdot,\mathbbm{1},\Delta,\epsilon,S,\ast,)$
is its dense $\ast$–Hopf algebra, then $H$ is the $\ast$–algebra generated by the smooth $\mathbb{C}$–valued functions
(28) $w_{ij}:G$ $\longrightarrow\mathbb{C}$
@@
with $A=(a_{ij})$. The coproduct of an element $g$ $\in$ $H$ is defined by
(29) $\Delta(g):G\times G\longmapsto\mathbb{C},\qquad\Delta(g)(A,B)=g(AB)\quad\mbox{ with }\quad A,\,B\,\in\,G.$
@@
(30) $\epsilon:H\longrightarrow\mathbb{C},\qquad g\longmapsto g(e),$
where $e$ $\in$ $H$ is the identity element; and
(31) $S(g):G\longrightarrow\mathbb{C},\qquad A\longmapsto S(g)(A)=g(A^{-1})$
for all $g$ $\in$ $H$.
Since $G$ is parallelizable, we have the identification $G\times\mathfrak{g}=TG=\mathfrak{g}\times G$, where $\mathfrak{g}$ is the Lie algebra of $G$. This diffeomorphism can be given by
(32) $(A,v)\longmapsto(d^{\prime}L_{A})_{e}(v),$
where $(d^{\prime}L_{A})_{e}$ is the usual de Rham differential $d^{\prime}$ at $e$ of the left product on $G$ with $A$; and by
(33) $(v,A)\longmapsto(d^{\prime}R_{A})_{e}(v),$
where $(d^{\prime}R_{A})_{e}$ is the de Rham differential $d^{\prime}$ at $e$ of the right product on $G$ with $A$.
Consider $d=d^{\prime}$. Under equation (32) we have $dg=g^{(1)}\otimes dg^{(2)}_{e}$ for all $g$ $\in$ $H$. By defining
@@
is a $\ast$–FODC over $\mathcal{G}$. Moreover, it is left–covariant under the map $\Phi_{\Gamma}=\Delta\otimes\mathrm{id}_{\mathfrak{h}}$.
On the other hand, under equation (33), we have $dg=dg^{(1)}_{e}\otimes g^{(2)}$ for all $g$ $\in$ $H$. Thus
$(\Gamma,d)\cong(\mathfrak{h}\otimes H,d)$
and it follows that $(\Gamma,d)$ is also right–covariant by the map ${{}_{\Gamma}}\Phi=\mathrm{id}_{\mathfrak{h}}\otimes\Delta.$ Hence, $(\Gamma,d)$ is bicovariant. According to [4, 6], there exists a right HH–ideal $\mathcal{R}\subseteq\mathrm{Ker}(\epsilon)$ such that
$\{\theta\in\Gamma\mid\Phi_{\Gamma}(\theta)=\mathbbm{1}\otimes\theta\}=\mathbbm{1}\otimes{\mathrm{Ker}(\epsilon)\over\mathcal{R}}={\mathrm{Ker}(\epsilon)\over\mathcal{R}}=:\mathfrak{qh}^{\#}$
@@
$\mathfrak{h}=\mathfrak{qh}^{\#}$
We claim that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon):=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,\mathrm{Ker}(\epsilon)\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$. In fact, consider the (restricted) quantum germs map
@@
Then, by equation (9), we obtain $\pi(g)=S(g^{(1)})dg^{(2)}=dg_{e}$. If $\mathfrak{g}^{\#}_{\mathbb{C}}$ is the complexification of the dual space of the Lie algebra $\mathfrak{g}$ of $G$, it is well–known that for $X=\{f\in C^{\infty}_{\mathbb{C}}(G)\mid\epsilon(f)=0\}$ we have
@@
with $X^{2}=\{\displaystyle\sum^{n}_{i=1}a_{i}\,b_{i}\mid a_{i},\,b_{i}\,\in\,X\,\mbox{ for some }\,n\,\in\,\mathbb{N}\}$ [7]. Moreover, for every $f$ $\in$ $X$,
@@
where $f]_{X^{2}}$ denotes the equivalence class of $f$ in $X/X^{2}$ [7]. In this way, for every $g$ $\in$ $\mathrm{Ker}(\epsilon)$ we get $dg_{e}=[g]_{\mathcal{R}}=[g]_{X^{2}}$. By this relation, $g$ $\in$ $\mathcal{R}$ if and only if $g$ $\in$ $X^{2}\cap\mathrm{Ker}(\epsilon)=\mathrm{Ker}^{2}(\epsilon),$ which proves that $\mathcal{R}=\mathrm{Ker}^{2}(\epsilon)$.
By definition, $\mathfrak{h}\subseteq\mathfrak{g}^{\#}_{\mathbb{C}}$. Since $\mathfrak{g}_{\mathbb{C}}\subseteq M_{k}(\mathbb{C})$, there exists a linear basis $\{v_{i}\}$ of $\mathfrak{h}_{\mathbb{C}}$ composed of linear combinations of the linear basis $\{E_{lj}\}$ of $M_{k}(\mathbb{C})$ ($E_{lj}$ is the matrix with $1$ in the $(l,j)$ position and 0 elsewhere). Then $(dw_{lj})_{e}(v_{s})$ corresponds to the $(l,j)$ position of the matrix $v_{s}$ . Consequently, there exist elements $\{(dg_{i})_{e}\}$, which are linear combinations of $(dw_{ij})_{e}$ $\in$ $\mathfrak{g}$, such that $(dg_{i})_{e}(v_{j})=\delta_{ij}$ with $\delta_{ij}$ being the Kronecker delta. Thus, $\mathfrak{g}^{\#}_{\mathbb{C}}\subseteq\mathfrak{h}$ and hence $\mathfrak{g}^{\#}_{\mathbb{C}}=\mathfrak{h}$. This implies that
(34) $(\Gamma,d)=(H\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d);$
so $(\Gamma,d)$ is a $\ast$–subFODC of the $\ast$–FODC
$(C^{\infty}_{\mathbb{C}}(G)\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d=d^{\prime})$
over $C^{\infty}_{\mathbb{C}}(G)$ of $\mathbb{C}$–valued differential 11–forms of $G$. Notice that it suffices to take convergent sequences $\{g_{i}\}^{\infty}_{i=1}$ $\subseteq$ $H$ in $(\Gamma,d)$ to recover $(C^{\infty}_{\mathbb{C}}(G)\otimes\mathfrak{g}^{\#}_{\mathbb{C}},d)$. It is worth mentioning that for the specific form $\Gamma=H\otimes\mathfrak{g}^{\#}_{\mathbb{C}}$, the right covariance map ${{}_{\Gamma}}\Phi$ is given by [6]
@@
for all $g$, $h$ $\in$ $H$.
@@
$d(hg)_{e}=\pi(hg)=\pi(h)\epsilon(g)+\epsilon(h)\pi(g)$
for all $h$, $g$ $\in$ $H$ and the right HH–module structure on $\mathfrak{g}^{\#}_{\mathbb{C}}$, given by
$\pi(h)\circ g=\pi(hg-\epsilon(h)g),$
simplifies to
(35) $\pi(h)\circ g=\epsilon(g)\pi(h).$
@@
of $(\Gamma,d)$. If g=abg=a\,b $\in$ $\mathrm{Ker}^{2}(\epsilon)$, we obtain
@@
$\mathfrak{g}^{\#\,\wedge}_{\mathbb{C}}=\otimes^{\bullet}\mathfrak{g}^{\#}_{\mathbb{C}}/A^{\wedge}=\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},$
where $\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}}$ denotes the exterior algebra of $\mathfrak{g}^{\#}_{\mathbb{C}}$. Thus (see equation (26))
(37) $(\Gamma^{\wedge}=H\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast).$
@@
(38) $(\Omega^{\bullet}_{\mathbb{C}}(G)=C^{\infty}_{\mathbb{C}}(G)\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast)$